OpenStudy (anonymous):

Please help Evaluate the polynomial function for the given values of the variable P(s)=2sᶟ-9s-10 P(1/2) and P(0)

OpenStudy (anonymous):

my pc froze had to restart website

OpenStudy (anonymous):

at 1/2 it is -57/4 at 0 it is -10

OpenStudy (anonymous):

need step by step explanations how did you get -57/4

OpenStudy (anonymous):

P(1/2) = 2(1/2)^3 - 9(1/2) - 10 P(1/2) = -57/4 P(0) = -10

OpenStudy (anonymous):

is that good enough?

OpenStudy (anonymous):

2(1/2)^3=1/4 9(1/2)=9/2 1/4-9/2-10 this is where i am stuck what do i do to get the correct answer

OpenStudy (anonymous):

okay so \[\frac{1}{4}-\frac{18}{4}- \frac{40}{4}= -\frac{1-18-40}{4} = \frac{57}{4}\]

OpenStudy (anonymous):

How do you do fractions chiarocuro?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

to add and subtract, you have to find a common denominator (preferably a LEAST common denominator) and change the fraction accordingly. Then add and subtract. to multiply, you just multiply numerator by numerator and denominator by denominator to divide, flip (the reciprocal) the fraction right after the dividing sign and then multiply

OpenStudy (anonymous):

lol you're silly. I mean how make fractions come out in equations? like you did there.

OpenStudy (anonymous):

lololols oops sorry in the equation editor type \frac{numer}{deno} replacing numer and deno for numbers cheers

OpenStudy (anonymous):

ok thanks. So is equations like latex? That's what I use in latex to do fractions.

OpenStudy (anonymous):

I'm never touched latex, so i couldn't tell you :x

OpenStudy (anonymous):

i need fractions for dummies p(1/4)=4(1/4)^3+9(1/4)-3 please explain this problem step by step i can not do fractions even the basic it nvr made sense to me

OpenStudy (anonymous):

\[4*\frac{1}{4}*\frac{1}{4}*\frac{1}{4} + \frac{9}{4} - 3\] \[\frac{1}{4}*\frac{1}{4} + \frac{9}{4} - 3\] \[\frac{1 * 1}{4 * 4} + \frac{9}{4} - 3\] \[\frac{1}{16} + \frac{9}{4} - 3\] \[\frac{1}{16} + \frac{9 * 4}{4 * 4} - 3\] \[\frac{1}{16} +\frac{36}{16} - 3\] \[\frac{37}{16} - \frac{3}{1}\] \[\frac{37}{16} - \frac{3 * 16}{1 * 16}\] \[\frac{37}{16} - \frac{48}{16}\] \[\frac{-11}{16}\] sorry that took so long...but that was a lot of equations :P

OpenStudy (anonymous):

mathlab says thats wrong its -83/16

OpenStudy (anonymous):

nope i used my calculator to check, i'm right. you can use matlab, but you don't know how to do fractions? please don't take offense! i'm just curious

OpenStudy (anonymous):

we do our homework in mathlab and i know how to do fractions just forgot exactly how its been so long

OpenStudy (anonymous):

copy this into matlab 4(1/4)^3+9(1/4)-3 ^-- thats the line you gave me it should give you -.6875, which is -11/16

OpenStudy (anonymous):

we might be talking about different things then nvm i'm talking about matlab you're talking about mathlab so nvm :P but i'm sure i'm right, as long as you gave me the right question but since you're trying to understand how to do fraction, just see if you can follow along the steps. because you want to know how to do it, not the answer

OpenStudy (anonymous):

right

OpenStudy (anonymous):

its a polynomial function given values of the variables p(-4) & p(1/4) P(r)=4rᶟ-9r-3 P(-4) and P(1/4)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

whoa...you still on this question? did you sleep?!?! O_O